Session: Oct 2014 (session 3)
Wavelength: 21 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (Jun Yang at Fri Oct 31 10:57:11 2014) Success:
The LO and BBC frequency setup was not correct until 19:45 UT. Because of the fixed bandwidth (250 - 500 MHz) of L-band filter, the LO frequency 1220 MHz in the vex file should be shifted down to 1000 MHz.

Svetloe (Zborovskiy A. at Fri Oct 31 12:07:51 2014) Success:

Badary (Zhukova, Sorokovikov at Fri Oct 31 12:08:47 2014) Success:
Pause for intensive geo session 20:06 to 22:00 UT (scans no0050 to no0074).

Zelenchukskaya (Bezborodov A. at Fri Oct 31 12:10:32 2014) Minor failures: ANT
Scans from no0012 to no0013 missed due antenna problems. Scans from no0044 to no0076 missed due to participation in intensive geo session.

Westerbork (Geert Kuper at Fri Oct 31 12:43:15 2014) Minor failures: SCHED
Scan 1,2,3,4 were byond Wb HA limits.

Torun (K. Borkowski at Mon Nov 03 06:44:10 2014) Success:
No known problems, although the operator noted some, apparently benign, error messages related to recording.

Effelsberg (Uwe Bach at Mon Nov 03 09:18:22 2014) Success:
Sources were below elevation limit until 16:31 UT.